A mobile is constructed of light rods, light strin
Answer #1

Here are the equations to balance.

d5 X 12 = d6 * m1

2.90 * 12 = 3.70 * m1

m1 = 9.405405405 gm

d3*m2 = d4*(m1+12)

m2 = 37.26126126 gm

d1*(12+ m1 + m2) = d2*m3
3.10 * (12 + 9.405405405 + 37.26126126) = 4.40 * m3

3.10 * 58.6666 = 4.40 * m3

m3 = 41.333333 gm